### v1

@@ 5.2.2 Experimental Results: Pruning Improvements
(a) The average degree of pruning achieved by different rule miners, as measured by the average proportion of lattice nodes evaluated. These measurements are plotted over increasing lattice sizes, showing that our Mono and Dual Rule Miners prune greater proportions of a lattice as its size increases.
(b) The average frequency of each rule type, measured as an average proportion of the lattice. These measurements are also plotted over increasing lattice sizes, showing that once all supporting sources are given to the LLM, valid retention and omission rules emerge in nearly half of the lattice nodes.
Figure 4: As more HotpotQA supporting sources are provided to a RAG system, our rule miners can prune larger portions of the lattice.
This trend is consistent with and supported by the patterns in the adjoining Figure 4(b). As more of the supporting sources are considered by the LLM, the number of omission rules decreases rapidly, but the number of retention rules increases rapidly. Once all supporting sources are included, then even when additional non-supporting sources are added, roughly half the nodes in the lattice are valid retention rules, and half the nodes are valid omission rules (these are not necessarily mutually exclusive). In short, Figure 4(b) indicates that there are generally fewer retention rules than omission rules, and Figure 4(a) illustrates how our algorithm therefore prunes more retention candidates.
Dual’s pruning proportion also increases as more supporting sources are considered, but until all three sources are included, pruning opportunities are negligible. This suggests that, so long as some supporting sources are not yet considered by the LLM, nearly all source subsets are promising candidates for at least one rule type. After three sources are added, Dual’s pruning proportion continues to increase at a linear rate.
5.2.3 Experimental Results: Mono vs. Dual
Figure 5: The average proportion of input subsets evaluated (as measured over two lattices) by Dual vs. two Mono runs (one for each rule type). Two Mono runs results in many overlapping evaluations (identical LLM inference calls); however, Dual’s caching mechanism allows it to avoid evaluating any subset more than once.
A truly naive Dual algorithm would make two naive Mono passes (one for each rule type), but since Figure 4(a) measures only the proportion of lattice nodes visited (i.e., at least once), it only expresses that Dual usually visits most nodes at least once. For this reason, we seek to compare the number of source subsets evaluated by a single Dual run to that of two Mono runs (one for each rule type). Figure 5 illustrates the results of this comparison, which are derived from the same HotpotQA RAG experiment. The results are quantified in terms of the proportion of evaluated subsets across two lattices, as Mono could evaluate each subset twice (once for each rule type). With caching enabled, our Dual algorithm will avoid evaluating any subset more than once, thereby preventing unnecessary LLM inference.
Figure 5 highlights how Dual evaluates far fewer total subsets than two Mono runs. For fewer sources, running Mono twice results in most subsets being needlessly evaluated twice. As more supporting sources are considered by the LLM, the total number of subsets evaluated over both Mono runs drops at a sublinear rate, which follows from the rule count trends in Figure 4(b). Meanwhile, the number of subsets evaluated by Dual decreases at a very slow rate, with nearly all subsets being evaluated until all three supporting sources are considered. Upon reaching six sources, the number of subsets evaluated under these separate strategies converges, suggesting that the degree of overlap in subsets implicated by opposing rule types generally decreases as more sources are considered.
5.3 Quantitative Evaluation: Synthetic Data
Finally, we repeat a similar experiment on a synthetic dataset. By doing so, we control for model nondeterminism and illustrate the limits of our algorithms’ pruning. In this experiment, we study a lattice corresponding to four elements (i.e., $|\mathbf{s}|=4$). We iterate every possible assignment of nodes’ validity, computing the number of valid rules that exist in each. For each of these validity assignments, we measure the number of lattice nodes that our Mono Rule Miner explores to find them. In short, we measure the degree of lattice exploration necessary to find every possible rule in every possible lattice of size $2^{4}$.
We first group these raw measurements by the number of valid rules, then average the number of visited nodes in each group. We plot the results in Figure 6(a). The red dashed line delineates the number of valid rules in the lattice; in order to find every rule, our algorithms cannot visit fewer nodes than this. The results indicate that, on average, our Mono rule miner terminates relatively soon after finding all valid rules. For the sizes tested, Mono visits only one or two additional nodes in most cases. We contend that this is reasonable in practice, given that a search needs to encounter some invalid rules in order to realize it can terminate early.
(a) Efficiency of Mono Rule Miner in a synthetic setting, seeking rules for each possible validity configuration (lattice) for an input set of size 4 (as described in Section 5.3).
(b) Efficiency of Mono Rule Miner in the HotpotQA setting, as measured over examples where six sources were provided to the LLM (as described in Section 5.2).
Figure 6: In both synthetic and real settings, the actual proportion of lattice nodes evaluated by our rule miners is very close to the theoretical minimum necessary to find all valid rules.
In Figure 6(b), we visualize similar statistics from the telemetry of our HotpotQA experiment. We focus only on LLM predictions that were made with six sources (i.e., $|\mathbf{s}|=6$), wherein there may exist anywhere from zero to $2^{6}$ valid rules of each type. Since there are relatively few data points, we present these data as a scatter plot. The plot shows that more valid omission rules are present than retention rules, consistent with Figure 4(b). More importantly, Figure 6(b), supports the conclusions evident from the synthetic data in Figure 6(a).

### v2

@@ 5.2.2 Experimental Results: Pruning Improvements
Figure 4: The average degree of pruning achieved by different rule miners, as measured by the average proportion of lattice nodes evaluated. These measurements are plotted over increasing lattice sizes, showing that our Mono and Dual rule miners prune more nodes as more distractor sources are considered (albeit at a slowing rate).
The results illustrated in Figure 4 demonstrate significant efficiency gains via pruning, avoiding many of the $2^{|\mathbf{s}|}$ nodes that a naive algorithm would evaluate. Generally, once all necessary sources are provided to the LLM (i.e., $|\mathbf{s}|\geq 2$), the rule miners deem additional sources inconsequential to the consistency of predicted answers and prune larger portions of the lattice. Figure 4(a) demonstrates that Mono Rule Miner can prune as much as 50% of the nodes when seeking retention rules in the HotpotQA setting, and as much as 70% when seeking omission rules. The synthetic setting results in Figure 4(b) reflect similar trends with slightly smaller rates, indicating clear asymptotic convergence at a 25% pruning rate for omission rules and 75% for retention rules. Across both settings, the proportion of this pruning increases with $|\mathbf{s}|$, albeit at a slowing rate.
Mono Rule Miner pruning rates vary significantly by rule type, but the trend flips in different settings: more omission rules are pruned than retention rules in the HotpotQA setting, but more retention rules are pruned in the synthetic setting. Naturally, the number of valid rules found mirrors this same pattern (see Figure 8 in Appendix B). In our RAG QA formulation, retention rules must identify source combinations sufficient to induce a consistent answer (e.g., combinations with both necessary sources), but omission rules need only identify a combination whose removal induces an inconsistent answer (e.g., any combination containing one or more necessary sources). Therefore, in a deterministic setting (like our synthetic setting), we expect more omission rules than retention, but the non-determinism introduced by using real LLMs with real data (like our HotpotQA setting) defies these expectations. Due to several factors discussed in Section 5.4.2, especially the fact that the LLM guesses consistent answers without all necessary sources, fewer omission rules hold than expected and more retention rules hold than expected.
(a) Mono vs. Dual Rule Miner lattice exploration as evaluated on a real LLM tasked with answering HotpotQA questions.
(b) Mono vs. Dual Rule Miner lattice exploration as evaluated in a synthetic setting that controls for stochasticity and consistency.
Figure 5: The average degree of pruning achieved by Dual Rule Miner as compared to two equivalent Mono passes (i.e., seeking retention rules, then omission). Plotted over increasing lattice sizes and w.r.t. 200% lattice exploration, Dual avoids reevaluating a significant proportion of lattice nodes, but to a lesser extent with additional distractor sources.
Figure 4 reports that, unsurprisingly, Dual Rule Miner evaluates most lattice nodes and evaluates more nodes than either Mono run individually. Importantly, it also confirms that Dual avoids evaluating any nodes more than once, a key advantage over the alternative strategy for finding both rule types: running Mono once for each rule type. Figure 5 provides a more accurate juxtaposition of pruning efficiency for Mono vs. Dual, rescaling “exploration” in terms of two lattice passes. Dual maintains an exploration rate near 100% in both HotpotQA and synthetic settings (decreasing and eventually converging at 75% in HotpotQA), while Mono evaluates nearly 200% (i.e., all nodes twice) in the worst case. As more distractor sources are considered, Mono’s exploration rate drops at a steady rate before eventually converging to Dual’s rate. In short, Dual is more efficient for smaller sets of sources.
5.2.3 Experimental Results: Scalability
Interpreting the same data under a different lens, Figure 6 visualizes and contrasts the scalability of each rule miner. Specifically, it plots the number of predictions made by each rule miner against the number of sources considered. Generally, all rule miners improve upon the naive algorithm that would evaluate all $2^{|\mathbf{s}|}$ lattice nodes. Moreover, the rate of scaling is consistent not only with the worst-case time complexity derived in Section 4, but is highly consistent across both the HotpotQA and synthetic settings.
By bounding this efficiency experiment to $1\leq|\mathbf{s}|\leq 10$, we also establish an upper bound on the number of sources reasonably supported by our rule miners at $|\mathbf{s}|=10$. We note that many real-world information retrieval (IR) systems (e.g., web search engines) seek the top-10 documents most relevant to users’ queries [JANSEN2006248, web-search-engines], making this an appropriate boundary for explaining real-world RAG LLMs that utilize such systems. Figure 6(a) shows that when $|\mathbf{s}|=10$ in the HotpotQA setting, Mono Rule Miner finds all omission and retention rules using approximately 250 and 500 predictions respectively, while Dual finds both using approximately 750. Therefore, when $|\mathbf{s}|=10$, the prediction count of Dual is similar to running Mono twice (minimal efficiency improvement), but as $|\mathbf{s}|$ decreases, Dual requires slightly less than double (moderate efficiency improvement). Given that many real-world RAG LLMs (e.g., LLMs tested in these experiments) can complete QA inference calls with sub-second response time, we assert that all rules can be found in a few minutes when $|\mathbf{s}|\leq 10$.
When $|\mathbf{s}|>10$, computing all rules can become intractable due to the exponential time complexity. This asymptotic boundary is well-studied in the association rule mining literature, and is largely unavoidable due to the exponential number of combinations that must be evaluated in the worst case (the entire lattice). The algorithms presented in this work adopt a design similar to the original Apriori algorithm [apriori], which ultimately navigates this exponential lattice level-order. More efficient successors to Apriori have emerged in the literature, and should inform successors of the present work, but the exponential time complexity remains. We point interested readers to works such as Eclat [eclat], whose depth-first traversal strategy improves I/O overhead, and FP-growth [fpgrowth], whose divide-and-conquer strategy avoids candidate generation.
In the next subsection, we present a case study that describes a method for tractably applying our rule miners when $|\mathbf{s}|>10$, so long as $k$ (the number of truly necessary sources) is a small constant (e.g., $k\leq 10)$. We contend that, although $k$ is not necessarily known upfront by users of RAG systems, it is unlikely that both $|\mathbf{s}|>10$ and $k>10$ in many practical settings. IR research has shown that in real-world search engines, the probability of a ranked document being relevant generally decreases monotonically as rank increases [analysis_rank_dist], suggesting that $k$ is likely small in RAG settings. Alternatively, rules can be mined for large $|\mathbf{s}|$ by partitioning them (i.e., find rules separately for smaller batches) or by merging them (i.e., find rules of coarser granularity).
(a) Rule miner predictions as evaluated on a real LLM tasked with answering HotpotQA questions.
(b) Rule miner predictions as evaluated in a synthetic setting that controls for stochasticity and consistency.
Figure 6: The average number of predictions made by different rule miners. These measurements are plotted over increasing lattice sizes, and are consistent with the exponential worst-case time complexity of our rule miners. Moreover, they establish $|\mathbf{s}|=10$ as a tractable upper bound for use of our rule miners with modern LLMs.
5.3 Scalable Red Teaming Case Study
We now present a case study focused on a large-scale, real-world scenario in the domain of AI safety. This case study follows an AI safety researcher whose team is developing advanced red teaming methods to safeguard a proprietary RAG LLM system. These methods probe RAG LLMs adversarially in an attempt to identify vulnerabilities, biases, and other harmful behaviors. The user is investigating a vulnerability recently uncovered in the proprietary LLM: malicious instructions can be injected into the RAG prompt through its retrieved sources. Given that users of the proprietary system are able to supply their own RAG sources, patching this vulnerability is imperative to ensure that the system’s safety training is not circumvented at inference time.
The user develops a novel solution that they believe will catch prompt injections and refuse to respond. To test the robustness of this solution, the team applies their AI safety expertise to craft 50 adversarial RAG sources with diverse injection strategies, all attempting to induce a response containing a specific “exploit hash”. Though the user can quickly test each adversarial source independently, they wish to test whether specific combinations of the adversarial sources are effective in exploiting the vulnerability. To efficiently identify any effective combinations, the user seeks rules that associate specific combinations with consistent, successful exploitation.
To generate rules, the user defines a simple output predicate function that checks the output for the exploit hash. The user then provides all 50 adversarial sources to Mono Rule Miner to find all retention rules, but encounters a problem. Given the large number of sources it is searching over, the rule miner is proceeding at a very slow rate and will never terminate. Thinking of ways to reduce the dimensions of the problem, the user realizes that since RAG ultimately concatenates all sources into a single prompt, the 50 sources can also be arbitrarily concatenated or grouped to reduce their count. Once rules are mined for coarse-granularity grouped sources, those not implicated by a rule can be discarded, and finer granularity rules can then be mined by dividing the remaining grouped sources again.
The user writes a script to automate these steps in a divide-and-conquer fashion. Beginning with 50 sources in one grouped source, Mono Rule Miner confirms that the exploit is induced (using 1 LLM inference call). The grouping is therefore divided into 2 grouped sources of size 25 and passed again to Mono, which finds a minimal rule implicating both grouped sources jointly (maximum of $2^{2}$ calls). Since both grouped sources contain one or more sources necessary to induce the exploit, the script continues by dividing each into two grouped sources of size 12 and 13, forming a new set of 4 grouped sources. Mono finds a single rule implicating the first and fourth grouped sources jointly (maximum of $2^{4}$ calls), allowing for the second and third to be discarded. The first and fourth are divided in two again and the algorithm continues similarly; each iteration runs Mono Rule Miner, discards grouped sources not implicated in minimal rules, then divides remaining grouped sources and recurses to solve them.
Once all grouped sources become indivisible (i.e., the original granularity is recovered), the algorithm terminates and returns said sources. After no more than 78 total LLM inference calls, the script provides the user with a single minimal rule at the original source-level granularity: retaining adversarial sources $\left\{\text{D8},\text{D44}\right\}$ consistently induces the exploit. This suggests that D8 and D44 are jointly but not independently effective in bypassing the user’s novel safety solution, perhaps due to interaction effects. With these new insights, the user has been made aware of the exploitative power of these sources when combined, and can study them to refine their solution.
5.4 Quantitative Evaluation: Rule Robustness
Building upon the real-world quantitative evaluation presented in Section 5.2, we present additional quantitative experiments that evaluate the robustness and quality of rules found by our rule miners. We first study the robustness of different “LLMs-under-test”, measuring the precision and recall of our rule miners in a real-world HotpotQA setting. Next, we study the robustness of the “LLM-as-a-judge” output predicate introduced in Section 5.2.1, using expert validation to estimate its precision and recall.
5.4.1 Experimental Setup
In this robustness evaluation, we follow and build directly upon the efficiency evaluation setup described in Section 5.2.1. The goal of the present evaluation is to measure the robustness of our rule miners (which mine “exact” rules) when used with real-world LLMs (whose non-determinism can induce answer stochasticity). We aim to measure and analyze the consistency of the “LLM-under-test” and “LLM-as-a-judge” components, since their stochasticity and propensity for error can lead to false positives and false negatives in their respective functions. Naturally, this evaluation focuses exclusively on the real-world setting, as the synthetic setting (introduced also in Section 5.2.1) deliberately controls for the aforementioned components. Therefore, we once again adopt HotpotQA to conduct a similar multi-hop QA experiment.
Building upon the previous HotpotQA experiment, we adopt the same definition for retention and omission rules, as well as the same implementation for output predicates $O_{consistent}$ and $O_{inconsistent}$. However, the shift in experimental focus from efficiency to robustness necessitates a different parameterization. We propose that the robustness of the LLM components can ultimately be quantified in terms of precision and recall by comparing actual vs. expected rules. After running a rule miner to obtain the “actual” rules, precision measures how many actual were expected (valid), while recall measures how many expected rules were actually output. We construct ground truth “expected” rules from the HotpotQA QA pairs to facilitate this computation; retention rules must retain both necessary sources to induce a consistent answer, while omission rules must omit one or more necessary sources to induce an inconsistent answer.
Using these metrics, we proceed similarly to the steps described in Section 5.2.1. We draw 50 QA pairs from HotpotQA training set at random, and fix each with exactly 5 sources: both necessary sources and three additional distractor sources. We pose these questions to three different commercial RAG LLMs, each initialized with a temperature of 1.0: OpenAI’s GPT-5 Mini (gpt-5-mini-2025-08-07), Google’s Gemini 2.5 Flash (gemini-2.5-flash), and Anthropic’s Claude Haiku 4.5 (claude-haiku-4-5-20251001). This enables us to contrast robustness across several commercial LLMs of similar capability.
We task Dual Rule Miner with finding all rules (of both types) for each LLM posed with each question, but do so repeatedly to obtain 10 sampled LLM answers for each. Due to their inherent stochasticity, LLMs may produce any number of different outputs when posed repeatedly with the same input prompt, potentially contradicting itself. Many works in both the machine learning and LLM literature have formalized inference strategies that embrace this reality (e.g., ensemble methods [ensemble_methods, llm_ensemble]), often improving overall accuracy by sampling multiple answers and computing a single final answer (e.g., via majority voting). Building upon these works, we compute a majority answer for each sample size in the range of 1 to 10, assessing the sample group as consistent only when 50% or more individual samples are judged consistent (see Appendix A for more detail). We then contrast and analyze the evolution of robustness over these increasing sample sizes, thereby measuring the robustness of our “exact” rules in the face of answer stochasticity.
5.4.2 Experimental Results: “LLM-under-test” Robustness
In Figure 7, we present the robustness of the different “LLMs-under-test”, reporting on four pairings: precision and recall over both retention and omission rules. In general, the chosen LLMs demonstrate similar robustness within each pairing, and across all samples sizes therein. With the exception of omission rule precision, for which all LLMs score approximately 100%, the results indicate that Gemini 2.5 Flash is the highest performing, followed closely by Claude Haiku 4.5 and then GPT-5 Mini. Despite using a moderate LLM temperature of 1.0 to permit stochasticity, all robustness measurements remain nearly constant as more samples are considered for the majority vote. This indicates that, despite the possibility of contradicting each other, the stochastic outputs of LLMs can (at least in similar QA settings) remain sufficiently consistent to produce exact rules, regardless of sample size.
In Figures 7(a) and 7(b), we observe that omission rules exhibit near-perfect precision but significantly lower recall. This behavior reflects the fact that omission rules are easier to satisfy; in theory, omitting a single necessary source will suffice to induce an inconsistent answer. Therefore, more omission rules should exist than retention rules, as established in Section 5.2.2. It follows that the omission rules found are almost always valid (high precision), since many exist. Consequently, there are even more omission rules that may not hold under a fallible LLM, resulting in some being missed (lower precision) in our “exact rule” formulation.
(a) Precision of omission rules generated.
(b) Recall of omission rules generated.
(c) Precision of retention rules generated.
(d) Recall of retention rules generated.
Figure 7: The precision and recall of rules generated using different “LLMs-under-test” over an increasing number of answer samples.
In Figures 7(c) and 7(d), we observe the opposite trend: retention rules exhibit near-perfect recall but significantly lower precision. Although all necessary sources should be required to derive the ground truth answer, we observe many instances where predicted answers are consistent nonetheless. We contend that this exceptional behavior stems from data quality issues in the HotpotQA dataset, as well as the LLMs’ unexpected propensity to leverage their pre-trained knowledge and “guess”. Consequently, many source combinations satisfy the retention rule criteria in practice despite omitting one or both of the necessary sources (low precision). Likewise, the presence of all necessary sources almost always induces a consistent answer, leading to the discovery of most expected retention rules (high recall).
Diving deeper to rationalize this unexpected behavior, we note that several flaws have been documented in HotpotQA questions, such as sources being mislabeled as non-relevant [ni-etal-2021-mitigating, thakur-etal-2025-hard] and the inadvertent introduction of reasoning shortcuts [jiang-bansal-2019-avoiding, chen-durrett-2019-understanding]. Qualitatively analyzing our experimental results, we find evidence that these flaws contributed to the validation of unexpected retention rules (false positives) and invalidation of expected omission rules (false negatives).
Specifically, we find that some distractor sources inadvertently contain information relevant to the question, and can be used to induce a consistent final answer without all necessary sources. For example, one question asks “Which rock group was formed earlier, Beastie Boys or The Shins?” There are two necessary sources, one for each band; each states the year the group was formed (1981 and 1996 respectively). However, some distractors are unexpectedly relevant by virtue of establishing pre-1996 release dates for some Beastie Boys songs. So long as “The Shins” necessary source is retained, a consistent answer can be inferred by also retaining any of these “release date” distractors.
Furthermore, we note that general HotpotQA data quality issues and inconsistencies may also affect our results. For instance, one question used in our experiments asks “When was the composer who wrote the music for Euryanthe born?”, but the ground truth answer “31 January 179719” is malformed. In this same example, the first necessary source establishes that “Carl Maria von Weber” wrote the music for “Euryanthe”, but the second necessary source states only the malformed birth date of the composer “Franz Peter Schubert”, who did not write “Euryanthe”. To illustrate another issue, we cite the question “Which profession does Boris Barnet have in common with Baltasar Kormákur?”, whose ground truth answer is “film director”. Despite the fact that HotpotQA questions should have a unique answer, we find that the two necessary sources indicate “actor” as another valid answer.
In cases where necessary information was truly absent, the LLM often chose to defy its instructions and “guess” the answer instead of strictly inferring it. Many of these guesses were consistent (and judged as such), perhaps due to the fact that the commercial LLMs tested likely trained on the same Wikipedia data from which HotpotQA was constructed. The remaining inconsistent guesses accounted for much of the stochasticity observed across answer samples. We note that answer diversity can be expected to increase as temperature is increased, and decrease as temperature is decreased. However, since diversity is offset by our “majority answer” sampling, we would not expect rule validity to be significantly impacted if our experiments were repeated with a temperature of 0.
5.4.3 Experimental Results: “LLM-as-a-judge” Robustness
To assess the reliability of the “LLM-as-a-judge” used across all of our quantitative experiments, we conduct an evaluation of its precision and recall utilizing ground truth answers. We randomly sample 100 judgments made by the “LLM-as-a-judge” during the robustness experiments (inspecting all 10 answer samples), drawn randomly across all QA pairs and “LLMs-under-test”. Each judgment is reviewed for correctness by an expert validator, who assesses its correctness by comparing the consistency of the predicted answer w.r.t. the HotpotQA ground truth answer. By tracking the number of examples that were accurately and inaccurately judged as “consistent” and “inconsistent”, we can compute the precision and recall of the “LLM-as-a-judge”.
Across the evaluated judgments, the “LLM-as-a-judge” achieves precision of 98.89% and recall of 100%, indicating that nearly all answers judged “consistent” were correct, and all truly “consistent” answers were correctly judged as such. The false positive rate is 0.90%, indicating that extremely few “inconsistent” answers were judged “consistent” by mistake. These results indicate that the judge is highly reliable at identifying answers consistent with the ground truth answer, and makes few mistakes in this QA setting. Moreover, we observe that the “LLM-as-a-judge” judges consistency at a semantic level, affording a level of nuance that makes it robust to unexpected inconsistencies.
Specifically, we observe that the “LLM-as-a-judge” is robust in identifying alternative answers that are consistent with the ground truth answer provided. For example, the HotpotQA question “Kevin Olekaibe plays for the team that competes in what league?” is labeled with the ground truth answer “NBA G League”. This league is known more generally as the NBA’s developmental league, and has been renamed multiple times (most recently to “NBA G League”, named after its current sponsor “Gatorade”). This question introduces ambiguity, as previous names for the league may or may not be consistent with ground truth. Expert validation finds that the “LLM-as-a-judge” exhibits semantic understanding of the answers in the context of the question, and judges accordingly. In one response to this question, the “LLM-as-a-judge” correctly inferred that “NBA Development League” is a more general term to describe the “NBA G League”, and correctly judged it as “consistent”.
Beyond robustness to diverse answers, we also observe robustness to diverse LLM responses. Our experiments show that Claude Haiku 4.5, despite being limited to 100 output tokens and prompted to output only the final answer, frequently outputs reasoning text that exhausts the allotted output tokens and is eventually truncated (i.e., before establishing a final answer). In most cases, these responses should be judged as inconsistent since a final answer was not clearly established, but reasoning sufficiently conclusive and aligned with ground truth could qualify as consistent. We note that across truncated responses observed via expert validation, those misjudged as “consistent” typically proposed the ground truth answer as a candidate answer, but were truncated shortly before articulating a final answer. These situations demonstrate how judgment of reasoning-heavy or partial answers is nuanced and potentially subjective, but the “LLM-as-a-judge” excels in this setting. Expert validation confirms that these truncated responses are rarely misjudged, as evidenced by the low false positive rate.